Bromine-82 has a half-life of 36 hours. A sample containing Br-82 was found to have an activity of

1.4 x 105 disintegrations/min. How many grams of Br-82 were present in the sample? Assume that

there were no other radioactive nuclides in the sample.


1
Expert's answer
2021-02-15T03:25:37-0500

As we know,

Decay constant, "k =\dfrac{0.693}{t_{\frac{1}{2}}}=\dfrac{0.693}{36\text{hour}}=0.01925hr^{−1}"

"N_o=1.4\times 10^5\text{disintegration/min}=\dfrac{1.4}{60}\times 10^5{\text{disintegration/hour}}"

Also,


"k=\dfrac{2.303}{t}log\dfrac{N_0}{N_t},"


"log\dfrac{N_0}{N_t}=\dfrac{kt}{2.303}=\dfrac{0.01925\times 36}{2.303}=0.3009"


or "\dfrac{N_0}{N_t}=Antilog0.3009=1.999=2"


"N_o=N_t\times{2}=\dfrac{1.4\times 10^5\times 2}{60}=4.66\times10^3\text{disintegration/hour}"


Grams of Br-82 present in the sample is "N_t-No=(4.66-2.33)\times 10^3=2.33\times 10^3"


="\dfrac{2.33}{4.66}\times 100=50" %


Grams of "Br-82=100-\dfrac{50}{100}(82)=82-41=41g"